2x squared + x -21=0
Found 2 solutions by nerdybill, KADINE:
Answer by nerdybill(7384)

Here's a site that explains a general method:
http://www.purplemath.com/modules/factquad2.htm
.
2x^2 + x -21=0
(2x+7)(x-3) = 0

Answer by KADINE(2)

2*^2+*_21=0(2*+7)(*_3=0